USE EXCEL SOFTWARE TO ANSWER ALL QUESTIONS SHOW YOUR WORK CLEARLY. PLEASE SUBMIT PRINTOUT FOR THIS HOME WORK. PLEASE DO NOT EMAIL HOMEWORK WITH EXCEL FILES
media%2F622%2F622fc2df-9551-416f-b19f-9a
Answer #1

Problem 3-3

Using compounding formula, we need to calculate the value of investment after 6 years, compounded monthly

A = P * (1 + r/12) ^ (n*12)

where, A = amount of money accumulated after 6*12 = 72 months = 3x

P = principal invested today = 1x

r = interest rate (to be derived)

n = no. of years; n*12 = 72 months

Therefore, 3 = 1 *(1 + r /12) ^ (6*12)

3 = 1 * (1 + r/12) ^ 72

Solving for r, using excel = 18.45% p.a.

Problem 3-4

Year 0 6 Cash Flow $ 20,000 5,000 -10,000 To calculate the value of X, use goal seek function in excel Value of X is such that, sum of discounted cash flow is0 Cash Flow $ PV Factor @ 6% Discounted Cash Flow Sum of Discounted Cash Flow $ 20,000 1.000 20,000 -5,000 0.943 -4,717 10,0009,054 0.890 | 8,900 0.705 |< PV factor for respective year, at 6% 6,383

Problem 3-5

Using compounding formula, we need to calculate the value of investment after 18 years, compounded annually

A = P * (1 + r) ^ n

where, A = amount of money accumulated after 18 years = $100,000

P = principal invested today = $20,000

r = interest rate (to be derived)

n = no. of years = 18 years

Therefore, $100,000 = $20,000 *(1 + r) ^ 18

5 = (1 + r) ^ 18

Solving for r, using excel = 9.35% p.a.